OpenStudy (anonymous):

T=2pi sqrt(m/k) {find m, if T=3, k=50}

OpenStudy (anonymous):

what's the problem ?

OpenStudy (anonymous):

solve for m

OpenStudy (anonymous):

why don't you simply use google calculator ?

OpenStudy (anonymous):

do take care of the units too.

OpenStudy (anonymous):

(T / 2pi) = sqrt (m/k) (T/2pi)^2 = m/k m = k (T/2 pi)^2

OpenStudy (anonymous):

yes.., 11.39